OpenStudy (anonymous):

Differential Equations Undetermined Coefficients y''(x)+y(x)=2^x

OpenStudy (anonymous):

would my trial solution be Yp=A^x?

OpenStudy (anonymous):

Y'p=xA^x-1

OpenStudy (anonymous):

I don't think what im doing is right....

OpenStudy (anonymous):

or A^x+B?

OpenStudy (anonymous):

Unfortunately I don't have much time to check into this problem, although it seems very interesting. What you're doing seems reasonable to me, I would also start with your trial solution to be: \[\Large y_p=A^x \] if that wont work try adding factors and constants to it, such that: \[\Large y_p=A^{cx}+(b) \] although I believe the constant wont do much, however make sure that you differentiate your trial right: \[\Large y=2^x=\left(e^{\ln 2}\right)^x=e^{ \ln 2 x} \]

OpenStudy (anonymous):

yeah I thought that adding constants wouldn't do much since when we derive they go away...

OpenStudy (anonymous):

So if you differentiate your trial you should get something like: \[\Large y=A^x=e^{\ln Ax} \] differentiation will lead to: \[\Large y'= \ln Ae^{\ln Ax}=\ln A \cdot A^x \] I believe this is where you went wrong in your differentiation.

OpenStudy (anonymous):

yeah that's what maple gave me ill try working it again

OpenStudy (anonymous):

sorry guys got physics class....thanks for help though!